### v2

# FlorDB: Multiversion Hindsight Logging for Continuous Training

Abstract.
Production Machine Learning involves continuous training: hosting multiple versions of models over time, often with many model versions running at once. When model performance does not meet expectations, Machine Learning Engineers (MLEs) debug issues by exploring and analyzing numerous prior versions of code and training data to identify root causes and mitigate problems. Traditional debugging and logging tools often fall short in managing this experimental, multi-version context. FlorDB introduces Multiversion Hindsight Logging, which allows engineers to use the most recent version’s logging statements to query past versions, even when older versions logged different data. Log statement propagation enables consistent injection of logging statements into past code versions, regardless of changes to the codebase. Once log statements are propagated across code versions, the remaining challenge in Multiversion Hindsight Logging is to efficiently replay the new log statements based on checkpoints from previous runs. Finally, a coherent user experience is required to help MLEs debug across all versions of code and data. To this end, FlorDB presents a unified relational model for efficient handling of historical queries, offering a comprehensive view of the log history to simplify the exploration of past code iterations. We present a performance evaluation on diverse benchmarks confirming its scalability and the ability to deliver real-time query responses, leveraging query-based filtering and checkpoint-based parallelism for efficient replay.
PVLDB Reference Format:
PVLDB, 17(1): XXX-XXX, 2024.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 17, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/ucbrise/flor.
1. Introduction
Model development for machine learning (ML) differs from traditional software engineering, focusing on experimental iteration rather than conventional debugging. The primary goal is to boost the predictive performance of models, leading developers to adopt empirical methods involving thorough logging and continuous updates to code and data (Xin et al., 2021). A recent study highlighted the importance of high-speed experimentation in this field: one participant stated, “the most important thing to do is achieve scary high experimentation velocity” (Shankar et al., 2024). This approach, however, generates a significant data management challenge, as engineers must handle numerous iterations of code, datasets, and logs.
For this aggressively agile approach to be successful, machine learning engineers (MLEs) must be able to retroactively examine past results to guide ongoing experiments (Garcia et al., 2018). One method to enable model developers to “query the past” is by generating and maintaining comprehensive experiment logs, but this is burdensome and unrealistic. Ensuring comprehensive logs requires (a) foresight to add logging statements to the code for all relevant metrics, without knowing which ones are relevant in advance, and (b) data management discipline to manage the corpus of log files (and requisite metadata) that ensues.
In this paper, we explore a different approach: we show it is possible to query the past of machine learning experiments efficiently and on-demand, even in the absence of experiment logs. We achieve this automatically on the developer’s behalf, by managing model training checkpoints and the corresponding versions of code and data, and exposing experimental history—both what was actually logged and what could have been logged—as a virtual database that can be queried with familiar APIs like SQL or dataframe libraries.
1.1. Multiversion Hindsight Logging
Hindsight logging (Garcia et al., 2020) is a lightweight record-replay technique that allows MLEs to add logging statements to a long-running Python program after it executes, and then incrementally replay parts of the program very quickly to generate the log outputs that would have emerged had the statements been in the code originally. This technique encourages MLEs to optimistically and flexibly operate with a minimal initial logging scheme, typically restricted to loss and accuracy metrics during the training phase. Additional logging is added retroactively, when evidence of issues arises (e.g., in deployment) and further key information is required for debugging.
We realized the concept of hindsight logging in Flor, a record-replay system specifically designed for model training (Garcia et al., 2020). Flor embodies two salient features: low-overhead checkpointing, and low-latency replay from checkpoint. Flor checkpointing is adaptive and runs in the background, facilitating low-overhead checkpointing, and limiting the computational resources expended during model training. Simultaneously, the system ensures reduced latency during replay by leveraging memoization and parallelism through checkpoint-resume.
While hindsight logging is a useful core technology, it is insufficient on its own to address typical ML workflows, which invariably involve many training runs. MLEs, in their pursuit of high-velocity experimentation, often wish to revisit not just the most recent run of an experiment, but also prior runs that were performed using different versions of code and data. This presents a complex challenge that goes beyond the capabilities of traditional logging and debugging tools.
To tackle this challenge, in this paper we introduce multiversion hindsight logging, and a system called FlorDB that provides an efficient and easy-to-use solution. Multiversion hindsight logging is designed to track and manage multiple versions of ML experiments. In doing so, it provides a more comprehensive and effective solution than simple hindsight logging, ensuring that MLEs can readily look back through their iterative development processes, thereby better understanding and learning from their experimentation history. Achieving these goals requires overcoming a set of technical challenges described below.
1.2. Contributions
FlorDB, a Multiversion Hindsight Logging system, provides a SQL-like query interface for ease of use, treating every version run as a set of rows, and each logging statement as a “virtual column”. FlorDB includes automatic version control through Git and is designed to interoperate with multiple ML experiment management tools. In addressing the inherent challenges of implementing Multiversion Hindsight Logging, FlorDB makes the following key contributions:
  1. (1)
    Historical Query Manager using a Unified Relational Model: FlorDB presents MLEs with a simple metaphor of a relational view of queryable log results—whether the log statements were previously materialized, or need to be generated on demand via hindsight logging. The unified relational model allows users to issue ad-hoc queries using familiar SQL or pandas, thereby simplifying the process of exploring historical code executions .
  2. (2)
    Multiversion Hindsight Logging as Acquisitional Query Processing: FlorDB extends the concept of Acquisitional Query Processing (AQP) (Madden et al., 2005) by seamlessly integrating data acquisition into query execution through experiment replay. This framework differs from traditional systems that query static datasets. Selected experimental versions undergo sequential processing: logging statements are propagated to generate required data, followed by the partial or parallel execution of a modified training script.
  3. (3)
    Highly Accurate Time Estimation for Replay Query Refinement: FlorDB incorporates highly accurate cost prediction for estimating the runtime of on-demand replay queries. By leveraging runtime profiling statistics collected during the initial record phase, FlorDB provides highly accurate (error $\leq 5\%$) replay cost estimates. This precision enables developers to refine their queries before being bogged down by unnecessarily lengthy execution times.
FlorDB’s integrated features provide machine learning engineers with a flexible relational abstraction to capture and query the extended histories of their ML experiments. By harnessing the framework of FlorDB, MLEs can more quickly iterate from prior attempts to successful models.
2. Scenario: Catastrophic Forgetting
Figure 1. Average training losses and F1-scores for Alice’s object detection model over the last 6 months. The model undergoes continuous training, with batches of labeled data added approximately twice a month. These batch dumps result in temporary fluctuations in the loss. The overall trend is stable.
Figure 2. BDD100K dashcam images with bounding boxes. Top row contains sample images used by Alice for fine-tuning on roundabouts; bottom row corresponds to images for which the model fails to detect pedestrians.
Catastrophic forgetting is a pressing issue in machine learning and is especially problematic in production settings (Shankar et al., 2024). This phenomenon occurs when a model, upon being fine-tuned for a new task, loses its ability to perform well on a previously learned task. This creates a balancing act between updating the model for new data or tasks and maintaining its performance on older tasks.
2.1. Object Detection in Autonomous Vehicles
Imagine an MLE named Alice, responsible for the ongoing fine-tuning of an object detection model used in autonomous vehicles. Operating in a dynamic environment, Alice frequently updates the model with bi-weekly batches of labeled data. Her current focus is the specialized task of object detection within traffic roundabouts. Standard metrics like loss functions and F1-scores, shown in Figure 1, serve as her evaluation benchmarks.
However, Alice encounters an unexpected problem. A colleague, Bob, informs her of persistent failures in pedestrian detection. Upon examining a representative set of images (refer to the bottom row of Figure 2), she verifies the issue.
2.2. Responding to Performance Regressions
To combat this specific case of catastrophic forgetting in pedestrian detection, Alice employs hindsight logging for two key actions:
  • Model Rollback: Alice would like to add hindsight logging statements to calculate the F1 score for pedestrian detection across previous model versions. This would allow her to identify the latest point where the model still successfully detected pedestrians, enabling an informed rollback.
  • Alert Tuning: To prevent the issue from recurring, Alice wants to add alerting logic that checks the value of the pedestrian detector F1 score. However she needs to avoid “alert fatigue”, when too many spurious alerts are issued. To ensure that her altering is calibrated, she again utilizes hindsight logging to ensure the F1 thresholds for alerting is not triggered by valid runs.
Alice’s scenario underscores a significant limitation in current approaches to model versioning and logging. Given the expansive history of model versions, it is neither practical nor efficient for Alice to manually revisit past iterations of training to insert additional logging statements, replay training, and analyze the results. Multiversion hindsight logging provides the high performance replay mechanisms for retroactive logging statement replay with automated back-propagation of those statements to previous versions of code for a comprehensive evaluation. This enables Alice to retroactively analyze metrics (e.g. F1 scores specific to pedestrian detection) across all historical versions.
Figure 3. Alice’s PyTorch training with Flor API.
projid tstamp filename hidden lr batch_size seed f1_score f1_round f1_ped
roundabouts 2023-06-23 train.py 500 0.001 32 81 0.7022215 0.90000000 0.5192609
roundabouts 2023-06-24 train.py 500 0.001 32 63 0.7043473 0.88589064 0.5590010
roundabouts 2023-06-29 train.py 500 0.001 32 157 0.6912616 0.86693724 0.5482590
roundabouts 2023-06-30 train.py 500 0.001 32 42 0.6994197 0.89759576 0.5170792
roundabouts 2023-08-27 train.py 500 0.001 32 213 0.6982518 0.84088466 0.2392322
roundabouts 2023-08-28 train.py 500 0.001 32 12 0.6770945 0.90000000 0.2493409
roundabouts 2023-08-29 train.py 500 0.001 32 333 0.7026935 0.86367500 0.2493730
roundabouts 2023-08-30 train.py 500 0.001 32 99 0.6818689 0.88571969 0.2351695
roundabouts 2023-08-31 train.py 500 0.001 32 475 0.6995291 0.87202768 0.2285919
roundabouts 2023-09-01 train.py 500 0.001 32 198 0.6850775 0.88872464 0.2265961
roundabouts 2023-09-02 train.py 500 0.001 32 94 0.7200135 0.89775851 0.2485293
roundabouts 2023-09-03 train.py 500 0.001 32 37 0.7030796 0.86352228 0.2305369
roundabouts 2023-09-04 train.py 500 0.001 32 292 0.7084305 0.89572726 0.2373573
roundabouts 2023-09-05 train.py 500 0.001 32 70 0.7090070 0.87411934 0.2364624
roundabouts 2023-09-06 train.py 500 0.001 32 553 0.6865721 0.84508553 0.2418998
Figure 4. Alice’s flor.dataframe before and after hindsight logging (black and green text, respectively).
3. User Experience & API
In this section, we explore how Alice leverages FlorDB [^1] Available on PyPI, installable via pip install flordb for addressing performance regressions in pedestrian detection, notably without pre-logged F1-scores for pedestrians. FlorDB enables her to effectively identify and revert to a more efficient model version, and fine-tune alert thresholds. Our discussion underscores the FlorDB API’s role in simplifying the ML workflow, facilitating rapid iteration and in-depth analysis for users with varied expertise. We illustrate the API’s functionality in key activities such as logging, parameter management, and hindsight logging for model refinement. Key API features include:
  • flor.log(name: str, value: T) -> T
    Logs a given value with the specified name. Useful for tracking variables or parameters during an experiment run.
  • flor.arg(name: str, default: T) -> T
    Reads a value from the command line if provided; otherwise, uses the default value. During replay, retrieves historical values. Good for setting configuration parameters and allowing for quick adjustments via the command line.
  • flor.loop(name:str, vals:Iterable[T]) -> Iterable[T]
    Functions as a Python generator that maintains global state between iterations. Useful for “indexing” replay and coordinating checkpoints.
  • flor.checkpointing(**kwargs) -> ContextManager[None]
    Acts as a Python context manager that defines the set of objects, such as model or optimizer, to be checkpointed adaptively — based on the size of the checkpoint and the loop iteration time (Garcia et al., 2020) — at flor.loop iteration boundaries.
  • flor.dataframe(*args) -> pd.DataFrame
    Produces a Pandas DataFrame of FlorDB log information with a column corresponding to each argument in *args. The DataFrame also contains “dimension” columns, such as the project id, version id, timestamp, and so on. It is the default view in FlorDB for querying log data and selecting model checkpoints to load, using either Pandas or DataFrame-compatible SQL engines like DuckDB (Raasveldt and Mühleisen, 2019).
3.1. Running Experiments
Alice executes her model training via the following command that starts her train.py script with specific hyper-parameters:
  python train.py --kwargs lr=0.001 batch_size=32
Alice’s script (3) uses native Python functionality to run her ML experiment. Hyper-parameters such as learning rate and batch size are passed in using standard command-line arguments, which makes hindsight logging seamless and capable of interoperating with other tools typically used in the ML workflow.
3.1.1. Args & Logs
In the context of running experiments, Alice utilizes the Flor API to enhance configurability and traceability. The flor.arg function is instrumental in parameterizing her experiments, allowing her to define hyper-parameters such as hidden_size, learning_rate, batch_size, and seed either via the command-line or by utilizing default values specified within the script (lines 5-8 in 3). This ensures adaptability and reproducibility of her experiment’s configuration. In addition, flor.log is employed to record the experiment’s loss metrics and associated metadata at each optimizer step (line 20 in 3). The logging of the loss captures the dynamics of the model’s performance, enabling Alice to monitor the training process.
3.1.2. Checkpointing on Loop Boundaries
Alice’s code defines the set of objects to be checkpointed using flor.checkpointing (line 10); within that context, Alice calls flor.loop (lines 12-13) to use the checkpoints and control record-replay. After each iteration of the outermost loop, which here corresponds to the completion of an “epoch,” FlorDB assesses whether to checkpoint. This decision is based on a balance between the checkpoint’s size and the iteration’s duration, adhering to a strategy from prior work (Garcia et al., 2020). Checkpoints are created adaptively, approximately once per epoch.
3.1.3. Automatically commit changes to Git
At the end of the experiment, FlorDB writes a JSON logfile containing the execution’s sequence of log records to the experiment repository, and commits all changes. This is so future users of the repository can view historical logs and arguments, and better reproduce the experiments. By committing after every run, Flor ensures historical versions of code are later available for hindsight logging.
3.2. Experiment Analysis and Hindsight Logging
Alice reviews FlorDB’s flor.dataframe output, comparing F1-scores across datasets (see 4). After receiving insights from her colleague, Alice shifts her analysis focus to the pedestrian F1-scores. By editing train.py to include a new flor.log for f1_ped, Alice leverages hindsight logging to populate this metric retroactively. She invokes the replay feature via CLI as follows:
python -m flor replay f1_ped "tstamp >= ’2023-06-23’"
This command extends f1_ped logging to any past experiments dated June 23, 2023 or later, in preparation for subsequent analysis.
Upon initiating the replay command, Alice is shown a schedule of experiments slated for replay and must confirm to proceed. After receiving confirmation, Flor restores historical configurations using flor.arg. Notably, during this replay, Flor bypasses the main loop (lines 12-22 in 3) to directly load the model’s final state, as F1-scores are computed post-training (line 25 in 3).
Alice views the updated results with the following command:
cols = [’hidden’, ’lr’, ’batch_size’, ’seed’]
cols += [’f1_score’, ’f1_round’, ’f1_ped’]
flor.dataframe(*cols)
This command now generates a table including the f1_ped metric (green text in 4) for entries dated June 23, 2023, or later. This new data reveals a previously unnoticed performance regression in pedestrian classifications, as Bob had indicated (tstamp $\geq$ 08-27).
To further refine the analysis and calibrate alert thresholds, Alice updates the validation function to add a new flor.log statement as part of an assertion that pedestrian F1-scores never fall below some threshold (line 21 in 3). For retroactive application, she again employs FlorDB’s replay feature:
python -m flor replay f1_alert \
"tstamp >= ’2023-06-23’ and tstamp < ’2023-07-01’"
This process selectively executes each epoch’s validation logic (lines 21-22), avoiding retraining (lines 13-20) by loading final epoch states from checkpoints. FlorDB’s control via flor.loop facilitates this selective process: the outer loop cycles through epochs using checkpoints, while the inner loop is bypassed.
Figure 5. FlorDB architecture diagram with subsection headers in parentheses.
4. System Architecture
FlorDB maintains a cross-referenced relational database that interconnects code versions in git, large files in your preferred storage solution (e.g. an S3 bucket), and metadata stored in the database. The design of FlorDB aims to provide a unified platform for the management of code, checkpoints, logs, and their evolution over time (5). This section outlines the mechanisms FlorDB uses for executing experiments and replaying them for hindsight logging, highlighting how it stores, manages and tracks the data and metadata of model training.
4.1. Storage and Data Layout
As shown in the bottom of 5, FlorDB’s storage architecture consists of the following three units:
  • Git Repository: Central to version control, it captures the state of the working directory and JSON log files post-execution. This facilitates a comprehensive tracking of code versions and more transparent sharing of execution metadata, enhancing experiment reproducibility and analysis.
  • Relational Database: Following execution and logging to JSON, FlorDB unpacks and normalizes data from the JSON logs to populate the relational database. This process transforms the semi-structured JSON data into a structured format, suitable for efficient querying and analysis.
  • Object Storage: This component is responsible for storing large files and checkpoints, integrating with the database system to offer scalable and flexible storage solutions, whether locally or in the cloud.
This structured approach to storage ensures that FlorDB is well-equipped to manage the complexities of code versioning, data checkpointing, and metadata storage.
4.2. Relational Schema
The relational schema of FlorDB is a structured and normalized view over the JSON logs. As depicted in 6, the logs table is central to this schema, managing the data and relevant metadata of experiments. The schema for the logs table includes:
  • projid: Basename of the cwd or Git repo.
  • tstamp: Datetime marking when the run was started.
  • filename: Name of the source file producing the log entry.
  • ctx_id: An integer acting as a foreign key to associate each log entry with a related entry in the loops table.
  • value_name: text descriptor for the variable being logged.
  • value: The actual data logged, stored as text.
  • value_type: Integer classifying the type of data logged.
Complementing the logs table, the loops table tracks the context of the execution flow (the call stack) as follows:
  • ctx_id: Unique identifier for a specific loop context.
  • parent_ctx_id: References the parent context in nested loops.
  • loop_name: The name or identifier of the loop.
  • loop_entries: Number of entries or iterations within the loop.
  • loop_iteration: Specific iteration count within a loop.
The “unpack” process involves converting logging objects from JSON format into the relational structure of the database. Each JSON log record is parsed. The data is then mapped to the relevant columns in both the logs and loops tables: A null loop indicating a top-level log entry, the log record is mapped solely within the logs table; nested loop objects encapsulate the hierarchical structure of an experiment’s iterations, and the corresponding log record is reflected in both logs and loops tables.
pt
loops
pt
ctx_id: integer parent_ctx_id: integer loop_name: text loop_entries: integer loop_iteration: integer
pt
logs
pt
projid: text tstamp: datetime filename: text ctx_id: integer value_name: text value: text value_type: integer
pt
ts2vid
pt
ts_start: datetime vid: text ts_end: datetime
pt
git
pt
vid: text filename: text parent_vid: text contents: text
pt
obj_store
pt
projid: text tstamp: datetime filename: text ctx_id: integer value_name: text contents: blob
Figure 6. Data Model Diagram in Crow’s Foot Notation. Tables with a white background are basic; those in gray are virtual.
4.3. Flor DataFrame
The flor.dataframe presents the content of the database in a single-table view (7). The operations involved join and transform the relational schema (6) into the final view. Initially, the code joins the logs table with additional dimensions of data, such as the loops table, which are essential for capturing the full context of each experimental run. This is followed by a series of transformations that pivot the combined data, turning what were once values into column headers, to create a wide-format table where each unique loop and log entry becomes a distinct column in the DataFrame (e.g. 4).
4.3.1. Domain Mapping
We use the flor.dataframe as the default Flor view because of its easy to understand application-level semantics. Specifically:
  • Running Experiments normal-$\rightarrow$ Adding Rows: The execution of a new experiment results in the addition of rows to flor.dataframe. Each row in the DataFrame signifies a discrete iteration within an experiment, collating related data such as metrics, hyper-parameters, and state.
  • Adding Logging Statements normal-$\rightarrow$ Adding Virtual Columns: The inclusion of logging statements in the source code induces FlorDB to add virtual columns to the view produced by flor.dataframe. These virtual columns can then be referenced in subsequent queries like any other column in flor.dataframe.
  • Replay from Checkpoint normal-$\rightarrow$ Backfilling Nulls: Replaying an experiment from checkpoints corresponds to the backfilling of null values in the flor.dataframe view. This capability is critical to achieving the hindsight logging abstraction that FlorDB supports.
The ML developer can envision the flor.dataframe as a universal relational view over the columns they have selected, where the rows represent individual experiment runs or, in the presence of loops, iterations within those runs. The columns in the flor.dataframe view are potentially infinite; they can be defined post-hoc and subsequently populated using replay from a checkpoint. This fluidity in defining and back-filling data exemplifies the relational model’s capacity to adapt to an experimenter’s evolving analytical needs.
    def dataframe(conn, *args):
        dataframes = []
        loops = pd.read_sql(
          "SELECT * FROM loops", conn)
        for val_name in args:
            logs = pd.read_sql(f"""
              SELECT * FROM logs
              WHERE value_name = "{val_name}"
              """, conn)
            logs = logs.rename(
              columns={"value": value_name})
            # Unroll loop context
            while logs["ctx_id"].notna().any():
                # Iterate until fixpoint
                logs = pd.merge(
                  left=loops, right=logs,
                  how="inner", on=["ctx_id"])
                ln = logs["loop_name"].unique()
                logs = logs.drop(
                  columns=["loop_name",
                  "loop_entries"])
                logs = logs.rename(
                  columns={"loop_iteration": ln[0]})
                logs["ctx_id"] = logs["p_ctx_id"]
                logs = logs.drop(
                  columns=["p_ctx_id"])
            logs = logs.drop(columns=["ctx_id"])
            dataframes.append(logs)
        all_joined = reduce(
          outer_join_on_common_columns, dataframes)
        cols = [c for c in all_joined.columns
          if c not in args] + list(args)
        return all_joined[cols]
Figure 7. Python code implementing data reshaping for flor.dataframe. This figure illustrates the sequence of operations—merging, transforming, and pivoting—used to construct the experiment view within FlorDB.
5. Acquisitional Query Processing
FlorDB extends the concept of Acquisitional Query Processing (AQP) (Madden et al., 2005), which involves efficient data acquisition during query execution. AQP was originally presented in the context of live sensing; here we acquire data by re-running training code from checkpoints. The process begins at the query construction stage, where criteria for selecting experimental versions are established (5.1). Each selected version has logging statements propagated to it in order to generate the required data (5.3). Subsequently, a modified training script is executed with a --replay_flor flag, enabling replay query operators to perform partial or parallel model training replay (5.2.1). In this section, we describe multiversion hindsight logging via the sequence of mechanisms by which FlorDB integrates data acquisition into query execution. This also paves the way for comprehensive post-hoc analysis and continuous model refinement.
5.1. Query Parsing
The flor replay operation begins with query construction, a step that determines the scope and precision of the retrospective analysis. This stage is crucial for identifying the code versions to be replayed and for setting the level of detail in the logging process. The query is formulated using the following command structure:
 python -m flor replay [-h] VARS [where_clause]
The components of the query include:
  • VARS: This argument specifies the logging variables that the user wishes to generate during the replay. These variables correspond to the log statements that will be regenerated across different code versions.
  • where_clause: This optional argument acts as a filter to refine the replay’s scope. It can be used to constrain the replay to certain conditions, such as experiments conducted within a specific timeframe or parameter set. The where_clause is evaluated against a flor.dataframe prior to replay.
By judiciously crafting the query, users can ensure that the replay operation is both focused on the necessary aspects of the experiment and efficient in terms of computational resource usage.
5.2. Empirical Cost Prediction & Plan Execution
Machine Learning (ML) training scripts can be long-running, and even selective replay can be time-consuming. Therefore, ML engineers need to know how long a replay will take, to decide whether further query refinement is needed. Although query cost estimation is a notoriously hard problem (Selinger et al., 1979; Yang et al., [n.d.]), we can accurately estimate costs for replay executions. This is because FlorDB profiles runtime information during the record phase. When initiating a Flor replay, these runtime statistics are queried using the flor.dataframe API call. This allows us to accurately estimate the time required for replay, enabling users to tailor their queries for efficiency. We categorize the cost estimation into four levels: s
  1. (1)
    Prefix Scan: Executes statements before the main loop, logging preliminary setup operations (e.g., lines 1-10 in 3. This level is useful when examining data preparation and featurization.
    • Cost Estimate: sum(prefix)
    • Subprocess: python train.py --replay_flor 0=-1
  2. (2)
    Suffix Scan: Executes the initial setup code (as in prefix scan), loads the end state of the outermost loop from a checkpoint, and runs the final script segments (lines 23-EOF in 3). This level provides a more detailed view, useful when examining the results of training (e.g. accuracy, recall, F1-scores).
    • Cost Estimate: sum(prefix + ckpt_load + suffix)
    • Subprocess: python train.py --replay_flor 0=0
  3. (3)
    Validation Scan. Steps through the main training loop to execute model validation logic (lines 21-22 in 3)., loading checkpoints once per epoch but skipping the nested training loop. This level is useful for examining the validation process in detail.
    • Estimate: sum(prefix + N*(ckpt_load + val_margin))
      where N is the number of epochs.
    • Subprocess: python train.py --replay_flor 0=1 1=0
  4. (4)
    Range Scan:Runs the training loop in-depth (lines 12-22 in 3) over a selected range to capture nested logging statements — used for the most granular analyses (e.g. of gradients). This level can be used individually or as part of a full replay in parallel, where each GPU is assigned a range of work.
    • Estimate: sum(prefix + ckpt_load + (k/N)*outer)
      where outer is the profiled time to run all N epochs, and k is the length of the selected range.
    • Subprocess: Let SEG be a list of int values denoting a contiguous range of epochs $i..j$ s.t. $0<i<j\leq N$; then,
      python train.py --replay_flor 0=\$SEG 1=1
5.2.1. Plan Execution
Once the replay’s depth level has been determined by a static analysis of the training script, and the end-user has accepted the replay plan, a replay subprocess is invoked for each version in the plan, with --replay_flor arguments formatted as strings and flags guiding Flor’s operations. As shown in the list above, Flor replay arguments can be used to break out of the execution in a prefix scan upon encountering the first flor.loop (--replay_flor 0=-1 where 00 is the index of the first flor.loop); or set to skip the first flor.loop loading its end checkpoint as part of a suffix scan (--replay_flor 0=0); or set to step into the the first flor.loop (--replay_flor 0=1), and so on. As shown above, validation scans and range scans control access to the outermost loop (index 00 or "epoch") and the nested loop (index $1$ or "step"), a key difference being that in validation scan the nested loop is skipped (1=0) but not in the range scan (1=1).
Table 1. Computer vision and NLP benchmarks used in our evaluation.
Model Model Size Data Data Size Objective Evaluation Application
ResNet-152 (He et al., 2016) 242 MB ImageNet-1k (Russakovsky et al., 2015) 156 GB image classification accuracy computer vision
BERT (Devlin et al., 2018) 440 MB Wikipedia (Foundation, [n.d.]) 40.8 GB masked language modeling accuracy natural language processing
GPT-2 (Radford et al., 2019) 548 MB Wikipedia (Foundation, [n.d.]) 40.8 GB text generation perplexity natural language processing
LayoutLMv3 (Huang et al., 2022) 501 MB FUNSD (Guillaume Jaume, 2019) 36 MB form understanding F1-score document intelligence
DETR (Carion et al., 2020) 167 MB CPPE-5 (Dagli and Shaikh, 2021) 234 MB object detection $\mu$-precision computer vision
TAPAS (Herzig et al., 2020) 443 MB WTQ (Pasupat and Liang, 2015) 429 MB table question answering accuracy document intelligence
(a) ResNet-152
(b) BERT
(c) GPT-2
(d) LayoutLMV3
(e) DETR
(f) Tapas
Figure 8. Comparative visualization of processing times for different Flor record-replay operations against the number of experiment versions. Note the logarithmic scale on the y-axis, emphasizing the time disparities between different operations.
5.3. Logging Statement Propagation
After selecting the versions and log variables for replay, but before the execution of individual Flor replay subprocesses, FlorDB adapts the code from the Git repository to incorporate new log statements. This involves propagating log statements from a newer version $Y$ to an older version $X$ ($X<Y$). The core challenge here is one of code-block alignment: identifying the most appropriate line in version $X$ to isert a logging statement that was originally in line $L$ of version $Y$. Code alignment is a fundamental task in software engineering, essential for understanding and managing changes in source files (Horwitz, 1990; Ladd, 1994). It allows developers to track modifications and maintain consistency across different versions of a file. In this context, GumTree emerges as a state-of-the-art technique for code block alignment.
5.3.1. Why GumTree is Sufficient for Our Use Case.
Code alignment, crucial for tracking and managing changes, uses GumTree (Falleri et al., 2014) for aligning code blocks, aiding in the accurate insertion of log statements across versions. The flor.loop() function establishes stable anchor points, with main and nested training loops serving as consistent markers across versions. This consistency, enhanced by GumTree’s alignment capabilities, ensures accurate log statement propagation. While GumTree is capable of aligning code blocks without Flor API anchors, the presence of these static loop identifiers all but ensures its efficiency and accuracy: in all instances of logging statement propagation evaluated in 6, this combined approach was successful.
To assess the effectiveness of these techniques, we conducted an evaluation focusing on accuracy and comparing GumTree with the Myers algorith (Myers, 1986; Falleri et al., 2014). The evaluation included synthetic benchmarks, randomly mutated PyTorch programs, and ecological benchmarks from GitHub. Our findings, detailed in a technical report (Dandamudi, 2021), show GumTree’s superior performance, especially in handling code refactoring and variable renaming.
6. Evaluation
This evaluation aims to demonstrate the key strengths of FlorDB: its ability to efficiently handle multiversion hindsight logging and its potential for significant speed-ups during replay. We showcase FlorDB’s versatility by applying it to a diverse range of model architectures, as detailed in Table 1. This table summarizes the models’ characteristics, evaluation metrics, and real-world applications. Building upon our prior work that established Flor’s low-overhead recording and parallel replay capabilities for single versions (Garcia et al., 2020), this study focuses on FlorDB’s ability to generalize and scale effectively across multiple versions. Our evaluation delves into two key areas:
  1. (1)
    Scalability: We assess FlorDB’s efficiency in handling growing numbers of model versions, ensuring linear scaling without performance degradation or cross-version interference.
  2. (2)
    Performance: We assess FlorDB’s techniques for accelerating hindsight logging replay. This includes accurate cost estimation for query refinement, capitalizing on cross-version parallelism for speed-ups from scale-out, and efficient utilization of within-version parallelism via checkpoint-based training resume.
Figure 9. Storage Requirements
6.1. Efficient Linear Scaling across Versions
Multiversion hindsight logging requires more computation than prior work on a single version, so we wish to demonstrate the scalability of FlorDB, and the absence of cross-version interference. We find that FlorDB scales efficiently and linearly with the number of versions.
6.1.1. Generalizing to $N$ Versions
We tested FlorDB across various model architectures, as shown in 8. Results show that the time complexity of FlorDB grows linearly with the number of versions (y-axis in log scale). This is a significant result, indicating the absence of interaction effects or accumulating overhead, even when dealing with big models or large datasets.
6.1.2. Interactive Response Times
A standout observation is the consistency in replay times across the different models. Despite the variability in the prefix phase due to variations between data preparation and featurization pipelines, most hindsight logging queries can be answered in near-real time (see brown lines in 8): even for more time-intensive record phases, there are multiple expedient replay options available to users.
6.1.3. Storage Requirements
9 shows the Git repository, despite frequent commits for dozens of versions, remains remarkably compact, under 5MB. However, checkpoints for large models stored on the shared drive can easily exceed hundreds of gigabytes (brown bars in the left pane). To address this, we propose “Kahneman Sampling” inspired by Kahneman (2011). This approach stores only two checkpoints per version: the “peak” checkpoint based on a chosen metric and the “end” checkpoint at training completion (olivegreen bars in the left pane). This minimal set surprisingly supports all replay-prefix and replay-suffix queries and enables restoring other checkpoints in half the time with 2x parallelism.
Figure 10. Boxplot distribution of errors between actual and expected execution times across different components of a replay execution. Components exhibit low errors on average, and estimates are highly correlated with projected runtime.
Figure 11. Comparison of runtime (top) vs rental cost (bot).
6.2. Speed-up through Selectivity & Scale-out
Estimating a replay query’s runtime is crucial for dynamic ML workloads where time-to-insight is paramount. This helps developers avoid prohibitively long execution times and efficiently utilize resources. As described in 5.2 and depicted in 10, we leverage record profiling statistics to provide highly accurate replay cost estimations (error $\leq 5\%$). These estimates empower developers to refine their queries until further filtering is no longer desirable. At that point, parallelism comes into play. Our prior work on Flor capitalizes on the embarrassingly parallel nature of training replay from checkpoints (Garcia et al., 2020). By scaling out operations across epochs (and also versions), Flor significantly speeds up replay. Therefore, our approach to reducing time-to-insight is two-pronged: first, accurate cost estimation for informed query refinement, and then, when further refinement is no longer desirable, efficient leverage of Flor’s checkpoint-resume parallelism for speed-up through scale-out.
6.2.1. Near-Ideal Parallelism
11 (top) depicts the scale-out ceiling, which represents the time, in minutes, required to run a single epoch of each experiment when parallelism is maximized: that is, when the degree of parallelism is set to equal the number of epochs. Prior studies confirm that Flor approaches ideal parallelism on a single version (Garcia et al., 2020), and FlorDB further leverages embarrassing parallelism from replay across versions.
6.2.2. Cost and Performance Considerations
Parallelism can accelerate computation, but it is also important to evaluate the associated costs. The estimated cost per epoch for various experiments on both CPU and GPU platforms is shown in 11 (bottom). Estimated cost is based on AWS EC2 instance rental fees, with and without a GPU, for the required execution time. While GPUs offer faster computation, they are associated with higher costs.
7. Related Work
ModelDB (Vartak et al., 2016) is a system designed to store, version, and manage ML models effectively. Similar systems include Weights & Biases (Biewald, 2020), VisTrails (Bavoil et al., 2005), and others (Vartak et al., 2018; Breck et al., 2019; Kumar et al., 2017). ModelDB allows users to log complete model metadata, including hyper-parameters, data splits, evaluation metrics, and the final model file. Moreover, it provides the ability to query over this metadata, making it an effective tool for analysis and comparison of different ML models and experiment runs. However, the focus of ModelDB is mainly on managing the straightforward metadata of deployed models, and it does not provide any features for hindsight logging. ModelDB’s focus on metadata is orthogonal to FlorDB’s multiversion hindsight logging facilities; the two systems could be used in concert or individually.
MLFlow (Zaharia et al., 2018) is an open-source platform that helps manage the end-to-end machine learning lifecycle, including experimentation, reproducibility, and deployment. It provides functionalities to log parameters, versioned code, metrics, and output artifacts from each run and later query them. Its modular design allows it to be used with any existing ML library and development process. Similar systems include TFX (Baylor et al., 2017), Airflow (Apa, 2016), Helix (Xin et al., [n.d.]), and others (Anil et al., 2020; Google Cloud, 2021). Like ModelDB, MLFlow has no support for hindsight logging or versioned log management. However, also like ModelDB, MLFlow and FlorDB can be fruitfully used together or separately.
FlorDB’s query model builds upon the notion of Acquisitional Query Processing (AQP) (Madden et al., 2005). Traditional query processing assumes that data is pre-stored and ready for querying; AQP introduced the concept of efficiently acquiring data as a part of the query process. The idea is particularly useful for applications like sensor networks, where querying can be expensive in terms of energy or computational resources. FlorDB’s multiversion hindsight logging also uses a form of AQP, where queries are not just made over pre-stored data, but also consider acquiring data through experiment replay. Unlike the original work on AQP, our acquisition task—i.e., multiversion hindsight logging—raises unique technical challenges of its own, forming the bulk of our work.
R3 (Li et al., 2023), with its record-replay-retroaction mechanism, is primarily focused on database queries and transactions, capturing and replaying states at the database level. This approach is efficient for debugging with low-overhead recording and storage-efficient replay. While R3 aims to enhance the debugging and testing capabilities of database-backed applications, FlorDB addresses the specific needs of machine learning model development and analysis with its hindsight logging and query processing features.
8. Conclusion
This paper introduces FlorDB, a system designed to address the unique challenges faced by machine learning engineers (MLEs) in managing the iterative model development process. FlorDB’s approach to multiversion hindsight logging, a record-replay technique, allows MLEs to add logging statements post-hoc, thereby enabling MLEs to “query the past.” FlorDB’s key contributions include a unified relational model for querying log results, automatic propagation of new logging statements across versions as part of acquisitional query processing, and a highly-accurate empirical cost predictor for replay query refinement. These features streamline the ML experimentation process, enabling more efficient analysis and faster iteration. We evaluate the system’s performance across various computer vision and NLP benchmarks, demonstrating the scalability of FlorDB when handling multiversion hindsight logging.